# Narrow line AGN selection in CEERS: spectroscopic selection, physical properties, X-ray and radio analysis

### v1

@@ 4.4 AGN prevalence
In Fig. 8, we show the fraction of the selected NLAGN among the CEERS MR sample. In particular, we consider as parent sample all the galaxies with a MR spectrum and with a secure redshift (217 sources). Then we divided the distribution into four redshift bins of between z=0 and z=9, and for each bin, we computed the fraction of AGN compared to the parent sample. The number of sources in the different redshift bins is 122, 85, 59 and 49, going from the lower to the higher redshift bin. We derived an almost constant fraction of AGN among the CEERS MR sources . We remind that this fraction do not represent intrinsic AGN fractions, given the fact the the original selection function of the spectroscopic survey is extremely hard to derive, but just a reference for the sample taken into account. This result agrees well with what was found for the NLAGN selection in the JADES survey by Scholtz et al. (2023a), where, however, the M1 and M2 and the R3O1 diagnostics diagrams were not involved in the AGN selection. Given the different sensitivities reached by the CEERS and JADES surveys, one would probably expect a lower fraction of NLAGN selected in the CEERS sample since, in most cases, the AGN selection is based on the detection of faint emission lines. At the same time, however, as we are going to show in Sec. 4.7, the median bolometric luminosity of the NLAGN in the CEERS sample is dex higher compared to those selected in JADES, and so the fact that the targets are brighter partially compensate for the shorter exposure times.
@@ 4.5 Velocity dispersion
In Fig. 9 we show the distribution of the intrinsic narrow line FWHM of the [O iii] 5007 emission line compared to the redshift of the sources. In particular, we considered only sources with the [O iii] 5007 line detected at S/N. To estimate the intrinsic FWHM, the observed FWHM retrieved by the fit was deconvolved by subtracting in quadrature the instrumental resolution of the grating at the observed wavelength of the [O iii] 5007 line. The instrumental velocity resolution was derived from the [point_source_lsf_f290lp_g395m_QD1_i185_j85.csv] line spread function file, which was calculated from the instrument model (Ferruit et al., 2022b), assuming a point-source geometry and a target located in the first MSA quadrant, at the centre of shutter (i,j)=(185,85); this procedure is described in de Graaff et al. (2024). The median value of the instrumental resolution is km/s.
The lower and upper boundaries of the intrinsic FWHM distribution are limited by the prior on the width of the line given to Qubespec, i.e. . None of the sources, after the subtraction of the instrumental resolution, shows a negative FWHM, but there is a non-negligible number of sources with an FWHMkm/s, meaning that in these cases we are really detecting emission lines at the limits of JWST resolution. Indeed, the cut in S/N and a careful visual inspection confirm beyond any doubt that these lines are real, and, actually, such narrow FWHM are expected given the trend of decreasing host galaxy stellar mass with redshift and given the fact that with JWST high-resolution spectrosocpy were recovered FWHM up to km/s (Maiolino et al., 2023). On the other hand, there are six sources with an intrinsic FWHM km/s. Four out of these six were classified as AGN using the diagnostic diagrams discussed above; in particular, three of these NLAGN are also low-redshift BLAGN reported in Fig. 17. Among the analyzed spectra, we did not find any significant residual from the fit of the narrow [O iii] 5007 line that could be indicative of the presence of an outflow component.
@@
The distribution of AGN and non-AGN, considering the errors, do not differ significantly from the other in any of the redshift bins. We also perform the Kolmogorov-Smirnov (KS) test on the global marginal distributions in FWHM of the two populations (histogram on the right panel), finding a p-value, which does not point towards different parent samples of the two distributions. This indicates that the NLAGN we spectroscopically selected among JWST spectra do not significantly impact their host-galaxy ISM, contrary to what was found in other studies on NLAGN samples, where, however, the NLAGN were selected based on other AGN activity tracers. For example, in the X-ray selected AGN sample of the KASHz survey, investigating sources at Cosmic Noon, (Harrison et al., 2016) found that of the targets have ionized gas velocities indicative of gas dominated by outflows and/or highly turbulent material, with [O iii] 5007 FWHMkm/s. We will further discuss the AGN impact on the sources studied in this work in the next sections.
@@ 4.6 Obscuration
In Fig. 10, we show the distribution of the values obtained from the Balmer line decrement. In particular, to derive the we considered the Small Magellanic Cloud (SMC) attenuation law (Gordon et al., 2003) for sources at and the Calzetti et al. (2000) attenuation law for sources at . While the Calzetti et al. (2000) attenuation law has been extensively used in the low- Universe to account for the effects of dust, the choice of the SMC attenuation law is more appropriate for the high- Universe, where galaxies are smaller and more compact than in the local Universe (Ono et al., 2023). In particular, we select only those sources that have both H and H detected. We assumed CASE B recombination, i.e. an intrinsic H /H ratio of 2.86444, ,(Osterbrock & Ferland, 2006). We also considered sources with upper limits in H to derive lower limits in the . The is computed in correspondence to the H emission line.
We note that a small number of galaxies in Fig. 10 scatter to either surprisingly high values of or else negative values, meaning an H /H line ratio less than the dust-free minimum value of 2.86. Even if sources with a dust-free H /H line ratio lower than CASE-B recombination have also been reported recently in the literature (Scarlata et al., 2024; Yanagisawa et al., 2024; McClymont et al., 2024), there is the possibility that a small fraction of these outliers are due to remaining systematics in the NIRSpec grating-to-grating flux calibration, i.e. when H and H are measured in different gratings. A similar result was indeed also observed in Shapley et al. (2023) on the same sample analyzed here. Even if present, these calibration issues do not affect our NLAGN selection, because the line ratios of the diagnostics presented above involve lines close to each other, and so generally in the same grating.
@@
We did not find a significant evolution of with redshift. We also note that the AGN population is characterized, in the lower redshift bin (), by a mag higher obscuration. At the average NLAGN obscuration is 1.24 mag compared to the 0.55 mag of non-AGN. On the contrary, when the whole NLAGN and non-AGN populations are taken into account, the average value of the obscuration across all redshifts is very similar, around 0.5 mag.
@@ 4.7 Bolometric luminosities
The way to compute the AGN bolometric luminosity is not unique, and it can be done starting from the X-ray luminosity, the luminosity of the BL emission or the UV continuum emission determined by the accretion disk. However, in our case, the majority of the sources are not X-ray detected, they do not have a BL emission, and their continuum is generally undetected. Therefore to estimate the bolometric luminosities, we rely on the dust-corrected narrow line fluxes of the H line, using the calibrations reported in Netzer (2009). These fiducial for the sample of NLAGN (with available H line) are reported with red squares in Fig. 11. In the same figure, we also report, with a fainter marker, the value of estimated from the same line fluxes but with the calibration used in Scholtz et al. (2023a) and taken from Hirschmann et al. (in prep). In this second case, the bolometric luminosity depends quadratically on the luminosity of the H line, while the Netzer (2009) calibration is linear, determining a difference in the estimated that can go up to dex (on average the bolometric luminosities computed using Hirschmann et al. calibrations are 0.8 dex lower). To verify the reliability of our fiducial values of we also considered the bolometric luminosity calibrations derived in Lamastra et al. (2009) where the NLAGN bolometric luminosities are derived from the dust corrected [O iii] 5007 line. The computed using the calibration in Netzer (2009) and those computed using the calibrations in Lamastra et al. (2009) agree well, with a median discrepancy of only dex.
@@
However, It is worth noting that all these calibrations assume that the lines used to derive are dominated by the AGN emission. This assumption is not necessarily true for the whole sample of NLAGN, and therefore, our should be generally taken as upper limits.
For the BLAGN at presented in Fig. 17, we computed the bolometric luminosities considering the same relation used in Harikane et al. (2023) to determine the bolometric luminosities of the BLAGN at of this sample, i.e the bolometric luminosity calibration derived in Greene & Ho (2005) from the broad H emission.
@@
In Fig. 11, we compare the distribution of of our sample with the AGN bolometric luminosities taken from the literature. In particular, we report AGN bolometric luminosities derived both from JWST spectroscopic studies and from pre-JWST surveys. Given the bolometric luminosities of AGN detected before the advent of JWST, it is clear that we are now able to sample a completely new regime in the luminosity-redshift space. Comparing the bolometric luminosities of our sample with those of the NLAGN selected sample from Scholtz et al. (2023a), we note that our targets are on average dex more luminous considering our fiducial calibration, but still more luminous even considering the luminosities derived with the calibration of Hirshman et al. (in prep). On the opposite, the AGN luminosities of our sample are comparable, in the same redshift range, with the distribution of derived from multiple samples of BLAGN detected with JWST.
4.8 Host Galaxies Properties
@@
In Fig. 12, we show the results from the SED-fitting performed on the whole sample of 217 sources as described in Sec. 2.4. In particular, we compare the trend with redshift of the ratio between the SFR derived from the CIGALE SED fitting and the one obtained from the main sequence (MS) reported in Popesso et al. (2023) at the stellar mass (, obtained from the fit) and redshift of each source. We note that sources at are generally distributed almost symmetrically with respect to the MS, with the median value of . On the contrary, at , sources appear to be systematically above the corresponding MS, with a median value of , i.e. SFRs 2-3 times larger than the respective MS SFR. Sources with larger SFRs are also the sources with the lower , some of them reaching . This means that the high- galaxies analyzed in this sample seem to have, on average, a higher MS normalization, being less massive and more star-forming than expected from the MS. Different works, taking advantage of JWST spectroscopy, already showed that high- galaxies are frequently in a state of intense or bursty SF (Dressler et al., 2023; Looser et al., 2023; Endsley et al., 2023), probably driven by a higher SF efficiency related to the lower metallicities of high- galaxies. However, another possibility that can determine this apparent bursty behaviour in our sample is a selection effect. Indeed, given that we are investigating a spectroscopic sample, we are probably more biased towards high SFR, in particular at high-z.
@@
For sources not selected as AGN, we also compared the SFR inferred from CIGALE to the values obtained from the H line (using the measured dust-corrected H luminosity and the relations reported in Shapley et al., 2023). On average, we find a good agreement, with with 0.25dex of scatter.
@@
The AGN distribution in Fig. 12 is in agreement with the distribution of the other sources not selected as AGN: considering the median distribution of the ratios of the two SFRs, even in different redshift bin, AGN and SFG are indistinguishable, as already found by other works (Ramasawmy et al., 2019). The AGN quenching effect on the host galaxy SF is still a largely debated topic (Bugiani et al., 2024; Man & Belli, 2018; Beckmann et al., 2017; Scholtz et al., 2018), and in this case, we do not note a negative impact of the AGN feedback on SF in the host galaxy.
@@
This finding is also in line with recent studies based on various cosmological simulations, according to which star formation quenching is not primarily driven by the instantaneous AGN activity but the integrated black hole accretion (as traced by black hole mass Piotrowska et al., 2022; Bluck et al., 2023; Scholtz et al., 2024). The AGN contribution to the final SED is measured in CIGALE by the parameter , which corresponds to the fraction of the AGN luminosity with respect to the galaxy one in the rest-frame 0.1-2 m. As expected, given the NLAGN nature of our sources, for most of the NLAGN, the fit returned a according to the scenario, already discussed in Sec. 4.5 and Sec. 4.6, that the AGN activity of the selected NLAGN is significantly buried by the host galaxy, in particular in the rest-frame optical part of the SED traced by the available photometry. This scenario, and the absence of a significant impact of the AGN activity over the host galaxy properties, is even more supported by the fact that we did not find relevant signatures of outflows in our spectroscopic analysis, in particular in the high- galaxies.
@@ 4.9 X-ray weakness
As described in Sec. 2.5 we performed a detailed analysis of the X-ray and radio counterparts of the selected NLAGN. Four out of five of the X-ray sources classified as AGN by the X-ray analysis in Buchner et al. (2015) are also classified as AGN by our NLAGN diagnostics. All the other 48 NLAGN do not show any X-ray counterpart. Therefore, we decided to perform an X-ray stacking analysis on our sample, using CSTACK as described in Sec 2.5. In particular, we decided to divide the sample of NLAGN in two redshift bins, equally populated, using as the dividing threshold. We also include in our analysis the BLAGN selected at in CEERS by Harikane et al. (2023), since none of them show any indication of X-ray emission and to perform a more complete analysis on the AGN sample among the CEERS MR spectra. We did not stack the sample of low- BLAGN presented in Appendix A because 3 out of 4 of these sources are X-ray detected (except CEERS-3129, which will be discussed later). The median redshifts of the three AGN samples are 2.37 for the low-redshift NLAGN, 5.50 for the high- NLAGN, and 5.43 for the high- BLAGN sample. In Fig. 13, we show the stacking maps in the SB and HB for the two NLAGN redshift bins and for the BLAGN, respectively. Given the median redshift of the samples, and considering the SB as the reference observing band since it provides the deepest constraint, we are investigating, on average, the following rest-frame X-ray energy bands: 1.7-6.5 keV for the low- NLAGN sample and keV for the other two samples. The stacking analysis did not reveal any X-ray detection in any of the sub-samples. This is quite surprising given the rest-frame energy range tested and the depth of the X-ray observations. To check whether the non-detections in the stacking of the NLAGN were due to possible galaxy contamination in the NLAGN sample, we also performed the same stacking but considering only NLAGN selected in more than two diagnostics diagrams, but again, the X-ray analysis did not reveal any detection.
Therefore, we wanted to test if such X-ray non-detections would be compatible or not with the expected X-ray luminosities of these objects.
@@
We used the luminosity relations derived in Duras et al. (2020) to compute the expected X-ray luminosities given the baseline dust-corrected bolometric luminosities computed in Sec. 4.7 for our sample of NLAGN. For the BLAGN, instead, we used the bolometric luminosities computed in Harikane et al. (2023) using the H broad emission. The comparison between the expected X-ray luminosities and the upper limits derived from the stacking is shown in Fig. 14. For all the AGN, the upper limits in their observed X-ray luminosities are largely below the expected X-ray luminosities, posing severe constraints on their X-ray production and emission. The X-ray deficit is dex for BLAGN and for the sample of NLAGN at , and is dex for the NLAGN at . The X-ray deficit of this last sample is particularly surprising, given the fact that most of these low- NLAGN were selected using the traditional BPT and VOT diagnostic diagrams, which are not expected to fail the NLAGN selection at low-.
@@
In Fig. 14 we also show the distribution of the X-ray detected sources in our MR spectroscopic sample together with the position of their X-ray stack.
In particular, for each X-ray source, we plot with a diamond the observed X-ray luminosity derived from their X-ray flux assuming , namely applying the same procedure as was done for the stack of the other sources. As expected, X-ray sources whose spectral analysis revealed larger column densities are on the left of the 1:1 relation, with a deviation up to dex for CEER-2989, whose spectral analysis revealed a column density of . This is because the spectral index , does not properly account for the harder spectrum of these heavily obscured sources. For each X-ray source, we also plot with circles the intrinsic, absorption corrected, X-ray luminosity derived in Buchner et al. (2015), which generally reduces the deviation from the 1:1 relation, especially if we take into account the large uncertainties on the intrinsic X-ray luminosities of the most obscured X-ray AGN. The same applies to the position of the stack of these X-ray sources: when we considered the observed X-ray luminosity (derived from the X-ray stacking) the point slightly deviates from the median expected X-ray luminosity, but if we correct for the absorption the X-ray luminosities (considering the median of the intrinsic X-ray luminosities on the x-axis), then the golden circle perfectly aligns with the median expected X-ray luminosity of these sources.
The same conclusions about the different samples can also be drawn by looking at the right panel in Fig. 14 where we compare the position of the sources with respect to the X-ray-bolometric luminosity relation of Duras et al. (2020). In the plot on the right we also show the lower limits derived in Maiolino et al. (2024a) from the X-ray stacking analysis performed on JWST selected BLAGN and NLAGN on the CDFS and Chandra Deep Field North (CDFN, Xue et al., 2016).
@@
Assuming that these sources are really AGN, the two possible scenarios that can justify such a lack of X-ray emission are:
Obscuration coming from large column densities, larger than the CTK limit ().
@@
A combination of the two possibilities above
Yue et al. (2024) analyzing a sample of so-called ‘Little Red Dots’ (LRD) showing clear AGN signatures (as the detection of a broad H component) and selected on various fields covered by JWST imaging and spectroscopy, found a similar result. The amount of the X-ray weakness of the LRD is comparable to what we found in this work for the sample of BLAGN at high-, while they did not investigate the population of NLAGN. In Yue et al. (2024), the authors suggest an intrinsic X-ray weakness as a possible origin of this discrepancy or a wrong AGN classification of these sources; in the latter case, the broad H component might be due to fast galactic outflows (with velocities km/s). In our case, however, we did not detect any signature of any outflow component in the forbidden lines of the BLAGN. More evidence and strong arguments against the outflow scenario have been presented by Juodžbalis et al. (2024a).
Also Maiolino et al. (2024a), considering a large sample of BLAGN and NLAGN spectroscopically selected in the JADES survey over the Chandra Deep Field South (CDFS, with 7Ms of Chandra observations Liu et al., 2017) show that almost all the AGN discovered by JWST lack any X-ray detection, even in the stack. In this work, the authors investigated different possible origins of this X-ray weakness, focusing in particular on the three possibilities reported above and finding indications in support of one or the other. In particular, they found an enhanced EW(H broad) in these sources that can suggest the Broad-line region to be distributed with a large covering factor around the central source acting as a dense dust-free absorbing medium, that can potentially reach Compton-thick column densities. On the other hand, the authors also find that many of the AGN newly discovered by JWST have features in common with the population of Narrow-line Seyfert 1 (NLS1), which are known to have a steep X-ray spectrum (and therefore more shifted towards a softer X-ray emission) and/or are characterized by high accretion rates, that are expected to result into a weaker X-ray emission, especially at high redshift, where higher rest frame energies are probed with Chandra.

### v3

@@ 4.4 AGN prevalence
In Fig. 8, we show the fraction of the selected NLAGN among the CEERS MR sample. In particular, we consider as parent sample all the galaxies with a MR spectrum and with a secure redshift (217 sources). Then we divided the distribution into four redshift bins of between and , and for each bin, we computed the fraction of AGN compared to the parent sample. The number of sources in the different redshift bins is 122, 85, 59, and 49, going from the lower to the higher redshift bin. We derived an almost constant fraction of AGN among the CEERS MR sources . This fraction does not necessarily represent the intrinsic AGN fraction at these redshifts, given the fact that the original selection function of the CEERS spectroscopic survey (necessary to correctly account for completeness corrections) is extremely hard to derive. However, this result agrees well with what was found for the NLAGN selection in the JADES survey by Scholtz23b. Given the different sensitivities reached by the CEERS and JADES surveys, one would probably expect a lower fraction of NLAGN selected in the CEERS sample since, in most cases, the AGN selection is based on the detection of faint emission lines. At the same time, as we are going to show in Sec. 4.7, the median bolometric luminosity of the NLAGN in the CEERS sample is dex higher compared to those selected in JADES, and so the fact that the targets are brighter partially compensate for the shorter exposure times.
@@ 4.5 Velocity dispersion
In Fig. 9, we show the distribution of the intrinsic narrow line FWHM of the [O iii] 5007 emission line compared to the redshift of the sources. In particular, we considered only sources with the [O iii] 5007 line detected at S/N. To estimate the intrinsic FWHM, the observed FWHM retrieved by the fit was deconvolved by subtracting in quadrature the instrumental resolution of the grating at the observed wavelength of the [O iii] 5007 line. The instrumental velocity resolution was derived from the [point_source_lsf_f290lp_g395m_QD1_i185_j85.csv] line spread function file, which was calculated from the instrument model (Ferruit22), assuming a point-source geometry and a target located in the first MSA quadrant, at the center of shutter (i,j)=(185,85); this procedure is described in deGraaff24. The median value of the instrumental resolution is km/s.
The lower and upper boundaries of the observed FWHM distribution are limited by the prior on the width of the line given to Qubespec; that is, . None of the sources, after the subtraction of the instrumental resolution, shows a negative FWHM, but there is a non-negligible number of sources with an FWHMkm/s, meaning that in these cases we are really detecting emission lines at the limits of JWST resolution. Indeed, the cut in S/N and a careful visual inspection confirm beyond any doubt that these lines are real, and, actually, such narrow FWHMs are expected given the trend of decreasing host galaxy stellar mass with redshift and given the fact that with JWST high-resolution spectroscopy were recovered FWHM up to km/s (Maiolino23c). On the other hand, there are six sources with an intrinsic FWHM km/s. Four out of these six were classified as AGN using the diagnostic diagrams discussed above; in particular, three of these NLAGN are also low-redshift BLAGN reported in Fig. 17. Among the analyzed spectra, we did not find any significant residual from the fit of the narrow [O iii] 5007 line that could be indicative of the presence of an outflow component.
@@
The distribution of AGN and non-AGN, considering the errors, do not differ significantly in any of the redshift bins. We also perform the Kolmogorov-Smirnov (KS) test on the global marginal distributions in FWHM of the two populations (histogram on the right panel), finding a p-value, which does not point toward different parent samples of the two distributions. This indicates that the NLAGN we spectroscopically selected among JWST spectra did not significantly impact their host-galaxy ISM, contrary to what was found in other studies on NLAGN samples, where, however, the NLAGN were selected based on other AGN activity tracers. For example, in the X-ray selected AGN sample of the KASHz survey, investigating sources at Cosmic Noon, (Harrison16) found that of the targets have ionized gas velocities indicative of gas dominated by outflows and/or highly turbulent material, with [O iii] 5007 FWHMkm/s. We further discuss the AGN impact on the sources studied in this work in the next sections.
@@ 4.6 Obscuration
In Fig. 10, we show the distribution of the values obtained from the Balmer line decrement. In particular, to derive the we considered the SMC attenuation law (Gordon03)555We used for the attenuation curve a fit to the average SMC-bar data points reported in Table 4 of Gordon03. We found that the analytic expression reported in their Eq. 4-5 does not correctly fit their observed data points at and has . () for sources at and the Calzetti00 attenuation law for sources at (with ). While the Calzetti00 attenuation law has been extensively used in the low- Universe to account for the effects of dust, the choice of the SMC attenuation law is more appropriate for the high- Universe (Reddy15; Shapley23), where galaxies are smaller and more compact than in the local Universe (Ono23). In particular, we select only those sources that have both H and H detected. We assumed CASE B recombination; that is, an intrinsic H /H ratio of 2.86666, (Osterbrock06). We also considered sources with upper limits in H to derive lower limits in .
We note that a small number of galaxies in Fig. 10 scatter to either surprisingly high values of or else negative values, meaning an H /H line ratio lower than the dust-free minimum value of 2.86. Sources with a dust-free H /H line ratio lower than CASE-B recombination have also been reported recently in the literature (Scarlata24; Yanagisawa24; McClymont24), but in our case these sources all have H /H line ratios compatible at 1-2 with 2.86. However, there is also the possibility that a small number of these outliers is due to small systematics in the NIRSpec grating-to-grating flux calibration; that is, when H and H are measured in different gratings. A similar result was indeed also observed in Shapley23 on the same sample analyzed here. Even if present, these calibration issues do not affect our NLAGN selection, because the line ratios of the diagnostics presented above involve lines close to each other, and so generally in the same grating.
@@
We did not find a significant evolution of with redshift. We also note that the AGN population is characterized, in the lower redshift bin (), by a mag higher obscuration. At the average NLAGN obscuration is 1.15 mag compared to the 0.5 mag of non-AGN. On the contrary, when the whole NLAGN and non-AGN populations are taken into account, the average value of the obscuration across all redshifts is very similar, around mag.
@@ 4.7 Bolometric luminosities
The way to compute the AGN bolometric luminosity is not unique, and it can be done starting from the X-ray luminosity, the luminosity of the BL emission or the UV continuum emission determined by the accretion disk. However, in our case, the majority of the sources are not X-ray detected, they do not have a BL emission, and their continuum is generally undetected. Therefore to estimate the bolometric luminosities, we rely on the dust-corrected narrow line fluxes of the H line, using the calibrations reported in Netzer09. These fiducial for the sample of NLAGN (with available H line) are reported with red squares in Fig. 11. In the same figure, we also report, with a fainter marker, the value of estimated from the same line fluxes but with the calibration used in Scholtz23b and taken from Hirschmann et al. (in prep). In this second case, the bolometric luminosity depends quadratically on the luminosity of the H line, while the Netzer09 calibration is linear, determining a difference in the estimated that can go up to dex (on average the bolometric luminosities computed using Hirschmann et al. calibrations are 0.8 dex lower). To verify the reliability of our fiducial values of we also considered the bolometric luminosity calibrations derived in Lamastra09 where the NLAGN bolometric luminosities are derived from the dust corrected [O iii] 5007 line. The computed using the calibration in Netzer09 and those computed using the calibrations in Lamastra09 agree well, with a median discrepancy of only dex.
@@
However, It is worth noting that all these calibrations assume that the lines used to derive are dominated by the AGN emission. This assumption is not necessarily true for the whole sample of NLAGN, and therefore, our should be generally taken as upper limits. For BLAGN at presented in Fig. 17, we computed the bolometric luminosities considering the same relation used in Harikane23 to determine the bolometric luminosities of the BLAGN at of this sample, i.e the bolometric luminosity calibration derived in Greene05 from the (dust corrected) broad H emission.
@@
In Fig. 11, we compare the distribution of of our sample with the AGN bolometric luminosities of other samples at comparable redshifts and taken from the literature. In particular, we report AGN bolometric luminosities derived both from JWST spectroscopic studies and from pre-JWST surveys. Given the bolometric luminosities of AGN detected before the advent of JWST, it is clear that we are now able to sample a completely new regime in the luminosity-redshift space. Comparing the bolometric luminosities of our sample with those of the NLAGN selected sample from Scholtz23b, we note that our targets are on average dex more luminous considering our fiducial calibration, but still more luminous even considering the luminosities derived with the calibration of Hirshman et al. (in prep). On the opposite, the AGN luminosities of our sample are comparable, in the same redshift range, with the distribution of derived from multiple samples of BLAGN detected with JWST.
4.8 Host galaxies’ properties
@@
In Fig. 12, we show the results from the SED-fitting performed on the whole sample of 217 sources that is described in Sec. 2.4. In particular, we compare the trend with redshift of the ratio between the SFR derived from the CIGALE SED fitting and the one obtained from the main sequence (MS) reported in Popesso23 at the stellar mass (, obtained from the fit) and redshift of each source. We note that sources at are generally distributed almost symmetrically with respect to the MS, with the median value of . On the contrary, at , sources appear to be systematically above the corresponding MS, with a median value of , in other words with SFRs 2-3 times larger than the respective MS SFR. Sources with larger SFRs are also the sources with the lower , some of them reaching . This means that the high- galaxies analyzed in this sample seem to have, on average, a higher MS normalization, being less massive and more star-forming than expected from the MS. Different works, taking advantage of JWST spectroscopy, have already shown that high- galaxies are frequently in a state of intense or bursty SF (Dressler23; Looser23; Endsley23), probably driven by a higher SF efficiency related to the lower metallicities of high- galaxies. However, this apparent above-MS behavior at high-z could also be due to a selection effect. Indeed, given that we are investigating a spectroscopic sample, we are probably more biased toward high SFR, in particular at high-z.
@@
For sources not selected as AGN, we also compared the SFR inferred from CIGALE to the values obtained from the H line (using the measured dust-corrected H luminosity and the relations reported in Shapley23). On average, we find a good agreement, with with 0.25dex of scatter.
@@
The AGN distribution in Fig. 12 is in agreement with the distribution of the other sources not selected as AGN: considering the median distribution of the ratios of the two SFRs, even in different redshift bin, AGN and SFG are indistinguishable, as already found by other works (Ramasawmy19). The AGN quenching effect on the host galaxy SF is still a largely debated topic (Bugiani24; Man18; Beckmann17; Scholtz18), and in this case, we do not note a negative impact of the AGN feedback on SF in the host galaxy.
@@
This finding is also in line with recent studies based on various cosmological simulations, according to which star formation quenching is not primarily driven by the instantaneous AGN activity but the integrated black hole accretion (as traced by black hole mass Piotrowska22; Bluck23; Scholtz24). The AGN contribution to the final SED is measured in CIGALE by the parameter , which corresponds to the fraction of the AGN luminosity with respect to the galaxy one in the rest-frame 0.1-2 m. As expected, given the NLAGN nature of our sources, for most of the NLAGN, the fit returned a according to the scenario, already discussed in Sec. 4.5 and Sec. 4.6, that the AGN activity of the selected NLAGN is significantly buried by the host galaxy, in particular in the rest-frame optical part of the SED traced by the available photometry. This scenario, and the absence of a significant impact of the AGN activity over the host galaxy properties, is even more supported by the fact that we did not find relevant signatures of outflows in our spectroscopic analysis, in particular in the high- galaxies.
@@ 4.9 X-ray weakness
As is described in Sec. 2.5, we performed a detailed analysis of the X-ray and radio counterparts of the selected NLAGN. Four out of five of the X-ray sources classified as AGNs by the X-ray analysis in buchner15 are also classified as AGNs by our NLAGN diagnostics. All the other 48 NLAGN do not show any X-ray counterpart. Therefore, we decided to perform an X-ray stacking analysis on our sample, using CSTACK as is described in Sec 2.5. In particular, we decided to divide the sample of NLAGN in two redshift bins, equally populated, using as the dividing threshold. We also include in our analysis the BLAGN selected at in CEERS by Harikane23, since none of them show any indication of X-ray emission and to perform a more complete analysis on the AGN sample among the CEERS MR spectra. We did not stack the sample of low- BLAGNs presented in Appendix A because 3 out of 4 of these sources are X-ray detected (except CEERS-3129, which will be discussed later). The median redshifts of the three AGN samples are 2.37 for the low-redshift NLAGNs, 5.50 for the high- NLAGNs, and 5.43 for the high- BLAGN sample. In Fig. 13, we show the stacking maps in the SB and HB for the two NLAGN redshift bins and for the BLAGNs, respectively. Given the median redshift of the samples, and considering the SB as the reference observing band since it provides the deepest constraint, we are investigating, on average, the following rest-frame X-ray energy bands: 1.7-6.5 keV for the low- NLAGN sample and keV for the other two samples. The stacking analysis did not reveal any X-ray detection in any of the sub-samples. This is quite surprising given the rest-frame energy range tested and the depth of the X-ray observations. To check whether the non-detections in the stacking of the NLAGN were due to possible galaxy contamination in the NLAGN sample, we also performed the same stacking but considering only NLAGN selected in more than two diagnostics diagrams, but again, the X-ray analysis did not reveal any detection. Therefore, we wanted to test if such X-ray non-detections would be compatible or not with the expected X-ray luminosities of these objects.
@@
We used the luminosity relations derived by duras20 to compute the expected X-ray luminosities given the baseline dust-corrected bolometric luminosities computed in Sec. 4.7 for our sample of NLAGN. For the BLAGN, instead, we used the bolometric luminosities computed in Harikane23 using the H broad emission. The comparison between the expected X-ray luminosities and the upper limits derived from the stacking is shown in Fig. 14. For all the AGNs, the upper limits of their observed X-ray luminosities are largely below the expected X-ray luminosities, posing severe constraints on their X-ray production and emission. The X-ray deficit is dex for BLAGNs and for the sample of NLAGNs at , and is dex for the NLAGNs at . The X-ray deficit of this last sample is particularly surprising, given the fact that most of these low- NLAGNs were selected using the traditional BPTs diagnostic diagrams, which are not expected to fail the NLAGN selection at low-.
@@
In Fig. 14, we also show the distribution of the X-ray detected sources in our MR spectroscopic sample together with the position of their X-ray stack.
In particular, for each X-ray source, we plot with a diamond the observed X-ray luminosity derived from their X-ray flux assuming , namely applying the same procedure as was done for the stack of the other sources. As expected, X-ray sources whose spectral analysis revealed larger column densities are on the left of the 1:1 relation, with a deviation up to dex for CEER-2989, whose spectral analysis revealed a column density of . This is because the spectral index, , does not properly account for the harder spectrum of these heavily obscured sources. For each X-ray source, we also plot with circles the intrinsic, absorption corrected, X-ray luminosity derived in buchner15, which generally reduces the deviation from the 1:1 relation, especially if we take into account the large uncertainties on the intrinsic X-ray luminosities of the most obscured X-ray AGN. The same applies to the position of the stack of these X-ray sources: when we considered the observed X-ray luminosity (derived from the X-ray stacking) the point slightly deviates from the median expected X-ray luminosity, but if we correct for the absorption the X-ray luminosities (considering the median intrinsic X-ray luminosity on the x axis), then the golden circle perfectly aligns with the median expected X-ray luminosity of these sources.
The same conclusions about the different samples can also be drawn by looking at the right panel in Fig. 14 where we compare the position of the sources with respect to the relation of duras20. In the plot on the right we also show the lower limits derived in Maiolino24_X from the X-ray stacking analysis performed on JWST selected BLAGN and NLAGN on the Chandra Deep Field South (CDFS, with 7Ms of Chandra observations liu17) and Chandra Deep Field North (CDFN, 2Ms xue16).
@@
s
Assuming that these sources are really AGNs, the two possible scenarios that can justify such a lack of X-ray emission are:
Obscuration coming from large column densities, larger than the CTK limit (), and/or
@@
Yue24, analyzing a sample of so-called “little red dots” (LRDs) showing clear AGN signatures (as the detection of a broad H component) and selected from various fields covered by JWST imaging and spectroscopy, found a similar result. The amount of X-ray weakness in the LRD is comparable to what we found in this work for the sample of BLAGN at high-, while they did not investigate the population of NLAGN. In Yue24, the authors suggest an intrinsic X-ray weakness as a possible origin of this discrepancy or a wrong AGN classification of these sources; in the latter case, the broad H component might be due to fast galactic outflows (with velocities km/s). In our case, however, we did not detect any signature of outflows in any of the forbidden lines of the BLAGN. More evidence and strong arguments against the outflow scenario have been presented by Juodzbalis24_rosetta.
Also Maiolino24_X, considering a large sample of BLAGNs and NLAGNs spectroscopically selected in the JADES survey over the CDFS and CDFN, show that almost all the AGNs discovered by JWST lack any X-ray detection, even in the stack. In this work, the authors investigated different possible origins of this X-ray weakness, focusing in particular on the three possibilities reported above and finding indications in support of one or the other. In particular, they found an enhanced EW(H broad) in these sources, which could suggest that the BLR is distributed with a large covering factor around the central source, acting as a dense dust-free absorbing medium that can potentially reach CTK column densities. On the other hand, the authors also find that many of the AGNs newly discovered by JWST have features in common with the population of narrow-line Seyfert 1 (NLS1), which are known to have a steep X-ray spectrum (with , and therefore more shifted toward a softer X-ray emission) and/or are characterized by high accretion rates, which are expected to result in a weaker X-ray emission, especially at high redshift, where higher rest frame energies are probed in the Chandra bands.